OpenStudy (anonymous):

What is the height of a cylinder if it has a radius of 3 centimeters and a volume of 254.34 cubic centimeters? (Use 3.14 for π.)

OpenStudy (anonymous):

V=B*h, Volume=Base*height here Base is an area for cylinders, the base is a circle you know V you know radius so you can find area of base, B with all that, you can find the height h easy

OpenStudy (anonymous):

\[A _{circle}=B=\pi r ^{2}\]